OpenStudy (anonymous):

What is the output value for the following function if the input value is 4.5? y = 2x + 0.5 2 4.5 9.5 25

OpenStudy (cwrw238):

replace x by 4.5 and work it out

OpenStudy (anonymous):

ouput = y input = x f(4.5) = 2(4.5)+0.5 = 9+0.5 = 9.5